OpenStudy (anonymous):

Describe the number of solutions for the equation. -3(x - 5) = 3x + 9

OpenStudy (anonymous):

-3 (x-5 ) = 3 ( x +3 ) -x +5 = x +3 2x = 3-5 x = -1 ie answer is single solution